OpenStudy (anonymous):

Write 2 log3 x + log3 5 as a single logarithmic expression.

OpenStudy (campbell_st):

well it uses 2 log laws 1st is the law of powers \[a \log(b) = \log(b^a)\] apply this law to the 1st term then 2, you need to law for multiplication \[\log(ab) = \log(a) + \log(b)\] hope it helps

OpenStudy (anonymous):

yes it does thx

OpenStudy (anonymous):

so, log3 x2 + log3 5 = log3(5x2)

OpenStudy (campbell_st):

that looks good.